#7. Let S (30) = 2+ - Find all numbers c that satisfy the conclusion of the Mean Value Theorem on the interval (1, 3).
Answer #1

f(x) = ara f (3) -f (1) 3+3 3- 1 - 1 - 9 = 6-10 = f(c)- 1+ - -2 = 1-9 9 = 3 + x²=3 a=t& Ja = [1,3]kindly give rating